The complete integral of √p+√q=3x


1
Expert's answer
2021-08-12T07:26:56-0400

Let f (p.q) = √p + √q – 3x = 0 …………….i

So p = ax + by + c ………..ii

f (a, b) = √a + √b – 3x = 0

So √b = 3x - √a

Or b = (3x - √a)^2

So p = 3x + (3 - √a)^2y + c